OpenStudy (anonymous):

Which equation is a point-slope form of the equation of this line? −2x+y+5=0 y=2x−5 (y+1)=−1 2 (x−2) (y−1)=2(x+2) Done Next © 2012 - 2015 PLATO Learning, Inc. All rights reserved. About

OpenStudy (mathmate):

The point-slope form of a line is the following: y-y1=m(x-x1) where m is the slope, and (x1,y1) is a point through which the line passes. The form y=mx+b is the slope-intercept, m=slope, b=y-intercept. For the given problem, you need to rearrange the given equation to the point-slope form and choose accordingly.